# A general framework for the generation of probabilistic socioeconomic scenarios: Quantification of national food security risk with application to the cases of Egypt and Ethiopia

### 07 Dec 2023 v1

@@ 3.3 A food security risk index within and across socio-economic scenarios
The setting above allows for the food security risk estimation within each one of the considered socioeconomic-climate probabilistic scenarios, with the term within meaning that we base our estimations on a single scenario out of the set ${\mathcal{Q}}$. However, it is not always feasible to identify with certitude the exact scenario that we are experiencing, or (as happens in the case of modelling within long time horizons) to know a priori that we will remain within one scenario due to the dynamic nature of the phenomenon. In such cases, one may wish to blend the various future perspectives in order to provide a risk assessment concerning the food security issue across scenarios. This is more realistic since in practice, he existence of a single universally accepted probability model (scenario) $Q$ is unusual. Instead of this, the set of probability models under different possible scenarios, ${\cal Q}$, provides different aspects of information concerning the evolution of the risk factors $Z$. This brings us to the realm of Knightian uncertainty Knight (1921), in which their is not a unique, universally acceptable probability measure for the description of the evolution of the risk factors affecting the phenomenon under study. This is a situation which requires a better way of providing the best estimate for the risk, than simply using the expectation ${\mathbb{E}}_{Q}[\Phi_{c}(Z)]$ under a unique probability measure. This would constitute what we call an estimator for the risk across scenarios; i.e. a measure of the risk of food security, under the uncertainty as to which scenario (equiv. model) will actually materialize. A proposal along these lines which is currently well accepted by the community of risk management, is within the context of convex risk measures and their robust representation (see e.g. Detlefsen and Scandolo (2005); Föllmer and Schied (2002); Frittelli and Gianin (2002)). Within this framework we calculate the risk through the variational representation
@@
Having adopted the fundamental conceptual framework of treating each scenario as a different probabilistic model (probability measure) for the risk factors $Z$ (mainly population factors in this study) we may now answer the question of robust estimation of the quantity of interest $I^{FS}_{c,t}$, for fixed $t$, using the convex risk measure of the form stated in (19). Following the suggestion in Papayiannis and Yannacopoulos (2018) and Petracou et al. (2022), in order to distinguish between the variou scenarios we adopt a metric space setting, according to which we consider the various scenarios in their natural setting, i.e., the metric space of probability measures, with a suitable metric $d(\cdot,\cdot)$, chosen so as to differentiate between the various probability measures in ${\cal Q}$, associated with different scenarios.
@@
4 Food security risk assessment for Egypt and Ethiopia under the SSP-RCP framework
@@ 4.1 Data and model assumptions
We follow the modeling approach presented in Section 3 for describing (a) the minimum caloric requirements (Section 3.1) and (b) food system capacity (Section 3.2) in order to provide estimates concerning food security risk for the cases of Egypt and Ethiopia through the risk indicator introduced in Section 3.3. For the estimation of the minimum calories intake component, it suffices to use the probabilistic estimates concerning the population structure evolution for Egypt and Ethiopia, since this quantity depends only on population (see the relevant discussion in Section 3.1). On the other hand, food system capacity at country level is described in terms of the two-layer model stated in 10-14, therefore relevant training data for the model calibration for both countries are required. For the model calibration task, historical data from the period 1990-2018 were used, since for this time period data are available or all quantities of interest that are taken into account in the modelling procedure. The required data for the model fitting procedure for Egypt and Ethiopia were mainly collected by the open database available from FAO, which includes historical data for food system capacity, food quantity imports and exports and land area used for agricultural purposes at national scale. Climate data and in particular, average temperature and average precipitation at national scale were obtained by the World Bank’s Climate database [^23] https://climateknowledgeportal.worldbank.org/ while water-stress records/estimations were obtained from the World Bank Database [^24] https://data.worldbank.org/. The variables that were used for the model calibration task and the sources from which the relevant data were acquired are illustrated in Table 8 in the Appendix.
@@
Concerning the model fitting procedure, Cobb-Douglas type parametric models were employed for the estimation of all the required parameters. For a full model description please see Section LABEL:app-4 in the Appendix. The fitting/estimation procedure has been performed separately for each country (not as panel data) using Ridge-type penalization to the typical OLS estimation approach, to treat potential instabilities caused by collinearity effects of the predictors. Note also that models are fitted separately to data and they are not fitted as a system of equations since different set of effects are considered for each quantity. The obtained model parameters for Egypt and Ethiopia are illustrated in Tables 9 and 10 in the Appendix with the relevant goodness-of-fit results. Concerning the food system capacity profiles, significant qualitative differences are not observed between the two countries. The technological and the year trend effects are at the same levels as far as the domestic food production is concerned. The only qualitative difference that is observed, concerns the effect of food exports where for Egypt the relevant coefficient displays a negative value, which possibly represents the difference in the food production capabilities and patterns between the two countries. Exported and imported quantity models seems to be quite similar for Egypt and Ethiopia. The food production mechanism of Egypt seems to be negatively related to potential temperature increases. Moreover, a negative dependence between the labour force occupied in food production activities in Ethiopia and GDP is observed, displaying potential social trends when income increases. However, most of the effects should be considered as aggregate effects and it is an interesting question for future research to see if these effects persist in finer scale models, e.g. involving a differentiation between food products or regions. However, such a task is subject to the availability of more detailed data, which are not currently available.
4.2 Food security risk estimation within SSP-RCPs
@@
For the projection tasks of this work, these two worlds are combined using the future estimates for GDP evolution as provided by MaGE under each one of the SSP scenarios, incorporating the conditional means for the population related quantities as obtained by the full probabilistic population model, keeping in mind the possible limitations and drawbacks from this approach. Climate drivers and in particular temperature and precipitation (in annual basis), are provided for each one of the scenarios in Table 3 from the Climate Change Knowledge Portal online database. Projections about the land area that will be occupied for agricultural activities, are performed under the assumption that no changes to the current land use policies are made. However, constraints as to the maximum area occupied for agricultural purposes according to each country capabilities in providing agricultural land, are applied. In this direction, for Egypt and Ethiopia we assign a maximum land area value that can be allocated to agriculture until 2050 with respect to the current capabilities of the countries. Therefore, to determine each country’s profile in agricultural land use, we estimated its trend for the last two decades, restricted by the upper bound that has been set by the physical limitations in land use. In this way, since land use depends on population, economy and climate factors (see model (14)), different estimates are obtained under each SSP-RCP scenario. Although more sophisticated models can be considered, taking into account thinner data about land use, under which the effect of land policies might be assessed by the proposed approach, we believe that such a task is beyond the scopes of this work. The key socio-economic and climate drivers which where used for the projections and the related models and sources that they were obtained from, are illustrated in Table 11 in Appendix C while elements concerning the uncertainty quantification of all main drivers that contribute to our modelling approach are provided in Appendix B.
@@
Combining the above projections we derive our estimations concerning food security under each scenario for Egypt and Ethiopia, in terms of the food security index introduced in Section 3.3 under four different perspectives (selection rules for the sensitivity parameter $\gamma$): (a) $\gamma=0$ (water stress is not taken into account), (b) very conservative perspective ($\gamma>0$ if $W_{c,t-1}>0.10$), (c) less conservative perspective ($\gamma>0$ if if $W_{c,t-1}>0.20$) and (d) no conservative perspective ($\gamma>0$ if if $W_{c,t-1}>0.40$). From the results in Figures 7, food security risk for Egypt seems to increase at the next decades, with higher rate at first years and with lower rate after 2040, under most of the SSP-RCP scenarios. Only for the SSP5-8.5 scenario after the year 2040 is indicated some small decrease on the risk. This picture is observed under all the different rules in selecting $\gamma$, except the case where the water risk component is omitted ($\gamma=0$). It is well known that Egypt is a highly water-stressed country and this picture is not expected to change in the near future. Without any doubt, availability of water resources is traditionally a critical issue for North Africa and especially Egypt, and it is still an unresolved problem. Therefore, the picture illustrated by the index for $\gamma=0$ is a quite unrealistic one, since the Egyptian agricultural sector strongly relies on water resources (see e.g. Christoforidou et al. (2023); Osman et al. (2016) and references therein) and therefore, food production activities are strongly related to water scarcity hazards mainly due to their current food production patterns, e.g. high emphasis in cereal production.
@@
For the case of Ethiopia, the estimated food security risk under no aggregation with the water risk ($\gamma=0$), displays similar behaviour (Figure 8). Under all SSP-RCP scenarios, the food security risk seems to decline at a very high rate, almost exponentially. However, this case is much more different than Egypt. Although Ethiopia is not considered as a water-stressed country (the relevant water stress indicator is below 40%), over the following years this situation is expected to change dramatically due to the strictly increasing trend in the historical water stress records. According to our model for water stress, by the end of 2050 the median water stress for Ethiopia is projected to lie in the interval between 300% and 500% (please see Figure 13 in Appendix C) under all the scenarios considered! This means that Ethiopia should be able to acquire an extra amount of water resources corresponding to 3-5 times of the amount of water resources that are currently used. Such an occasion will significantly affect the food security risk for the country and unavoidably will affect the food production activities. Adopting the water stress penalization perspectives introduced in Section 3.3, a more realistic picture concerning the food security risk evolution in the next years is obtained. In particular, even under the NC perspective the food security risk is expected to greatly increase at next years under all SSP-RCP scenarios due to the forthcoming water scarcity hazard that is predicted by the water stress model. The same picture is observed under all perspectives (NC, LC and VC) indicating very significant food security risk to the end of 2050 and especially even in the middle scenario SSP2-4.5.
4.3 Food security risk estimates across SSP-RCPs
Although it is useful to obtain the food security risk estimate within each SSP scenario, this risk evaluation is characterised by the main drawback that it is not robust to the uncertainty concerning to which scenario materialises. Since the policy maker needs to take into account all possible scenarios in order to make a robust decision, we employ the relevant robust risk assessment approach discussed in Section 3.3. In terms of a toy example, we provide three different types of policy makers (or experts [^26] Please see the very interesting paper Sutherland and Burgman (2015) concerning the task of incorporating experts opinions): (a) one of complete ignorance, so all possible outcomes are weighted equally, (b) an optimistic one, where higher probability to scenarios more favourable for the environment is placed and (c) a pessimistic one, where higher probability to scenarios that are less favourable for the environment is allocated. These types are illustrated in Table 6.
@@ C. Pessimistic 1/100 1/25 1/10 1/5 3/20 1/2
In Figure 9 estimates for the food security index under the NC, LC and VC selection rules for $\gamma$ are illustrated across the SSP-RCP scenarios. For Egypt, the estimations under all perspectives concerning water stress almost coincide, and a quite homogeneous tendency for all expert opinions is displayed, indicating an increasing trend of concave type in food security risk until the year 2040, and then a small decrease. Clearly, this is a consequence of the high hazard concerning water scarcity that Egypt is exposed to. For the case of Ethiopia, the food security risk estimates are also very homogeneous, indicating increasing trend for food security risk under all perspectives and types of experts, evolving at almost exponentially rate as time progresses. Therefore, this robust approach in aggregating scenarios indicates that both countries are about to face important food security risks, however, the situation in Ethiopia is predicted to get worse at much higher pace unless proactive measures are taken. In any case, the provided across scenarios estimates for both countries, certify their robustness providing very homogeneous predictions which seem to be quite insensitive to alternations of the weights (different types of experts) concerning the considered set of SSP-RCP scenarios.
@@ 5 Conclusions
5.1 Limitations and Extensions
The various models employed in this paper are unavoidably subject to various assumptions and limitations. These are:
  • Raftery’s population model upon which our probabilistic scenario modelling was based, does not explicitly take into account population control policies or climate effects, but only implicitly through the past data on which the Bayesian model is trained, which may affect the outcomes of the model. The same comment is true for MaGE. In our modelling approach we introduce climate effects exogenously, through the use of RCP scenarios. The explicit introduction of such effects in the above models is clearly beyond the scope of the present work, and is an interesting subject for future extensions.
  • The data used for calibrating the model for food capacity were from the period before 2018 and therefore may not account for potential shocks of phenomena arising, such as e.g. pandemics or political events that happened after that. This is unavoidable due to the fact that such events are rather recent resulting to the unavailability of sufficient data for the period 2019-2023, during which these events happen. However, using techniques from extreme value theory the effects of these shocks may be assessed by stress testing the current model.
  • Clearly, the proposed methodology could be extended to construct a more detailed food security model incorporating detailed information on the infrastructure of the local food production structures, dietary patterns, etc. However, such a model would require collection of detailed data which for many countries are currently unavailable, at least in a reliable form. Therefore we chose to construct a rather aggregate model and a resulting food security risk index that relies on publicly available and easily accessible and reliable data.
5.2 Concluding Remarks
@@
In this paper a general methodology for producing probabilistic socioeconomic scenarios compatible with the SSP-RCP framework and for assessing food security risk was proposed. The probabilistic scenarios represent the effects of the inherent uncertainty more efficiently than points estimates and are therefore better suited for projecting important socioeconomic quantities into the future. As an application of this methodology the issue of food security is considered, where a plausible index for its assessment is provided and a framework for the derivation of quantitative evaluations per scenario and future projections is proposed. While projections along scenarios are important, in practice there is uncertainty as to which scenario materializes. To address this question, a suitable framework for creating projections across scenarios using the concept of convex risk measures and their robust representation in the presence of model uncertainty is also proposed.
Our approach is illustrated by an application on assessing food security risk in two major countries in the upper Nile region, Egypt and Ethiopia. Examining each SSP-RCP scenario separately, our results show that under all possible trajectories, Egypt is not expected to face a serious food security risk for the next 30 years. On the contrary, although under all examined scenarios Ethiopia’s food security index shows an increasing trend, the pressure in country’s water resources is expected to rapidly and exponentially grow in the next years to levels that cannot be satisfied (about 3 to 5 times the country’s water resources would be needed till the year 2050). Such estimates should trigger the alarm for Ethiopia’s policy makers to design sustainable policies in terms of water use that should allow the country to avoid serious food security issues. The same conclusions carry over when we integrate all the above information, under three different hypothesis about the probability of scenario occurrence (i.e. complete ignorance, optimistic and pessimistic). Therefore, Ethiopian policy-makers should focus on possible pathways of food system transformation such as technology adoption, conflict reduction, strengthening resilience of the most vulnerable, lower the cost of nutritious foods, tackling poverty, etc. WHO et al. (2021) in an effort to mitigate the effects of the continuous food balance deficit, as well as of the sharpened food insecurity that is foreseen after 2030 due to rapidly increasing pressure in country’s water resources.

### 14 Dec 2023 v2

@@ 3.3 A food security risk index within and across socio-economic scenarios
The methodology described above facilitates the estimation of food security risk within each of the considered socioeconomic-climate probabilistic scenarios, where ‘within’ implies basing our estimations on a singular scenario from the set $\mathcal{Q}$. However, identifying with certainty the exact scenario being experienced, particularly over long time horizons, is often not feasible due to the dynamic nature of the phenomena involved. In such instances, it becomes pragmatic to integrate various prospective scenarios to assess food security risk ‘across’ scenarios. This approach reflects reality more closely, as the existence of a single, universally accepted probability model (scenario) $Q$ is rare. Instead, the set of probability models under different potential scenarios, $\mathcal{Q}$, offers diverse insights into the evolution of risk factors $Z$. This scenario leads us into the domain of Knightian uncertainty Knight (1921), where there is no singular, universally acceptable probability measure for describing the evolution of risk factors affecting the phenomenon under study. In such contexts, deriving the best estimate for the risk extends beyond merely computing the expectation $\mathbb{E}_{Q}[\Phi_{c}(Z)]$ under a unique probability measure. This necessitates an estimator for the risk ‘across’ scenarios – a measure of food security risk under the uncertainty of which scenario (equivalently, model) will actually materialize. One well-accepted approach in the risk management community within the context of convex risk measures and their robust representation involves using the variational representation (see e.g. Detlefsen and Scandolo (2005); Föllmer and Schied (2002); Frittelli and Gianin (2002)). In this framework, we calculate the risk through its variational representation.
@@
Having adopted the fundamental conceptual framework of treating each scenario as a different probabilistic model (probability measure) for the risk factors $Z$ (mainly population factors in this study) we may now answer the question of robust estimation of the quantity of interest $I^{FS}_{c,t}$, for fixed $t$, using the convex risk measure of the form stated in (19). Following the suggestion in Papayiannis and Yannacopoulos (2018) and Petracou et al. (2022), in order to distinguish between the various scenarios we adopt a metric space setting, according to which we consider the various scenarios in their natural setting, i.e., the metric space of probability measures, with a suitable metric $d(\cdot,\cdot)$, chosen so as to differentiate between the various probability measures in ${\cal Q}$, associated with different scenarios.
@@
4 Food Security Risk in Egypt and Ethiopia
@@ 4.1 Data and model assumptions
Adhering to the modeling approach presented in Section 3, for (a) the minimum caloric requirements (Section 3.1) and (b) the food system capacity (Section 3.2), in this section we estimate the food security risk in Egypt and Ethiopia, as per the risk indicator introduced in Section 3.3. For estimating the minimum calorie intake component, we primarily utilize probabilistic projections of population structure for both countries, as this quantity is predominantly dependent on population (refer to Section 3.1 for details). On the other hand, the food system capacity at the country level is delineated through the two-layer model described in 10-14, necessitating relevant historical data for model calibration. The calibration task employed data from the 1990-2018 period, during which comprehensive data for all relevant quantities were available. The data for model fitting for Egypt and Ethiopia were principally sourced from the FAO’s open database [^23] https://www.fao.org/faostat/en/#data, which provides historical information on food system capacity, food quantity imports and exports, and land area used for agricultural purposes at the national level. Climate data, specifically average temperature and precipitation, were obtained from the World Bank’s Climate database [^24] https://climateknowledgeportal.worldbank.org/, while water-stress records and estimations were sourced from the World Bank Database [^25] https://data.worldbank.org/. The variables utilized in the model calibration and the sources of the corresponding data are detailed in Table 8 in the Appendix C.
@@
For the model fitting procedure, we employed Cobb-Douglas type parametric models to estimate all necessary parameters. A comprehensive description of the full model is available in the Appendix C. The fitting and estimation process was conducted separately for each country (not as panel data), incorporating Ridge-type penalization into the typical OLS estimation method. This approach was chosen to address potential instabilities arising from the collinearity effects of the predictors. It is also important to note that models were fitted individually to data, rather than as a system of equations, due to the consideration of different sets of effects for each quantity. The resulting model parameters for Egypt and Ethiopia are presented in Tables 9 and 10 in the Appendix C, along with the corresponding goodness-of-fit results. In terms of the food system capacity profiles, there are no significant qualitative differences observed between the two countries. The technological effects and the year trend are comparable for domestic food production. A notable qualitative distinction is seen in the effect of food exports for Egypt, where the relevant coefficient is negative, possibly reflecting differences in food production capabilities and patterns between the two countries. Models for exported and imported quantities appear quite similar for both Egypt and Ethiopia. Notably, Egypt’s food production mechanism shows a negative correlation with potential temperature increases. In Ethiopia, a negative relationship is observed between the labor force in food production activities and GDP, hinting at potential social trends with increasing income. However, these effects should be interpreted as aggregate influences, and it would be an intriguing area for future research to examine if these effects persist in more granular models, such as those differentiating between food products or regions. Such an endeavor, however, hinges on the availability of more detailed data, which is currently not accessible.
4.2 Food security risk within SSP-RCPs
@@
For the projection tasks of this work, these two worlds are combined using the future estimates for GDP evolution as provided by MaGE under each one of the SSP scenarios, incorporating the conditional means for the population related quantities as obtained by the full probabilistic population model, keeping in mind the possible limitations and drawbacks from this approach. Climate drivers and in particular temperature and precipitation (in annual basis), are provided for each one of the scenarios in Table 3 from the Climate Change Knowledge Portal online database. Projections about the land area that will be occupied for agricultural activities, are performed under the assumption that no changes to the current land use policies are made. However, constraints as to the maximum area occupied for agricultural purposes according to each country capabilities in providing agricultural land, are applied. In this direction, for Egypt and Ethiopia we assign a maximum land area value that can be allocated to agriculture until 2050 with respect to the current capabilities of the countries. Therefore, to determine each country’s profile in agricultural land use, we estimated its trend for the last two decades, restricted by the upper bound that has been set by the physical limitations in land use. In this way, since land use depends on population, economy and climate factors (see model (14)), different estimates are obtained under each SSP-RCP scenario. Although more sophisticated models can be considered, taking into account thinner data about land use, under which the effect of land policies might be assessed by the proposed approach, we believe that such a task is beyond the scopes of this work. The key socio-economic and climate drivers which where used for the projections and the related models and sources that they were obtained from, are illustrated in Table 11 in Appendix C.
After modeling the minimum food requirements and food system capacity, we proceed to produce estimates for both quantities for the period 2019-2050 for Egypt and Ethiopia under each SSP-RCP scenario listed in Table 3. Population projections under each SSP scenario, as presented in Section 2.2, are readily available from the probabilistic model. For other socioeconomic indicators, several macroeconomic models, including the MaGE model (discussed in Section 2.2 and Fouré et al. (2013); Fontagné et al. (2022)), provide projections based on these population forecasts under the SSP scenarios. The MaGE model, utilizing data from UN and IIASA [^27] https://iiasa.ac.at/, offers global projections up to 2100 for socioeconomic activities across all countries for each SSP scenario.
Although it is theoretically possible to generate individual trajectories for socioeconomic quantities using MaGE for each population trajectory within scenarios, this approach would be computationally intensive and potentially conflict with MaGE’s focus on pointwise estimates, necessitating a new macroeconomic model. In this study, we generate projections for socioeconomic variables that align with the probabilistic population projections by substituting the standard population inputs in MaGE with the conditional means for population under each SSP, derived from the simulated trajectories appropriately assigned to each SSP scenario as described in Section 2.2. This approach provides estimates for GDP and labor that are consistent with the SSP probabilistic modeling philosophy. However, limitations and drawbacks of this approach are acknowledged.
Climate drivers, specifically annual temperature and precipitation, for each scenario in Table 3, are sourced from the Climate Change Knowledge Portal online database. Projections for agricultural land use assume no changes in current land policies, with constraints applied based on each country’s capacity for agricultural land provision. For Egypt and Ethiopia, we set a maximum agricultural land area up to 2050, estimating its trend over the last two decades within physical land use limitations. This results in different agricultural land use estimates under each SSP-RCP scenario, based on population, economy, and climate factors (see model (14)). While more intricate models could be used to analyze detailed land use data, this is outside the scope of this study. The key socio-economic and climate drivers used for these projections, along with the models and sources from which the data were obtained, are detailed in Table 11 in the Appendix C.
@@
Integrating the projections discussed previously, we obtain estimates for food security under each SSP-RCP scenario for Egypt and Ethiopia, using the food security index outlined in Section 3.3. These estimates are derived under four different perspectives for setting the sensitivity parameter $\gamma$: (a) $\gamma=0$ (excluding water stress), (b) a VC perspective ($\gamma>0$ if $W_{c,t-1}>0.10$), (c) a LC perspective ($\gamma>0$ if $W_{c,t-1}>0.20$), and (d) a NC perspective ($\gamma>0$ if $W_{c,t-1}>0.40$). According to the findings in Figures 7, the food security risk for Egypt is projected to increase in the coming decades, with a higher rate initially and a slower rate after 2040 under most SSP-RCP scenarios. A notable exception is a slight decrease in risk post-2040 under the SSP5-8.5 scenario. This trend is consistent across all methods of selecting $\gamma$, except when the water risk component is omitted ($\gamma=0$). Given Egypt’s status as a highly water-stressed country—a situation unlikely to change soon—the exclusion of water risk ($\gamma=0$) yields an unrealistic portrayal. Water resource availability has long been a critical issue in North Africa, particularly in Egypt, and remains unresolved. Consequently, the depiction by the index for $\gamma=0$ does not align with reality, considering the strong dependency of the Egyptian agricultural sector on water resources (see, for example, Christoforidou et al. (2023); Osman et al. (2016) and references therein), and the significant impact of water scarcity hazards on food production, notably in cereal production.
@@
For Ethiopia, the estimated food security risk, when water stress is not aggregated into the calculation ($\gamma=0$), shows a similar pattern (see Figure 8). Under all SSP-RCP scenarios, the risk appears to decrease markedly, almost exponentially. However, Ethiopia’s situation differs significantly from Egypt’s. Despite currently not being classified as a water-stressed country (with the water stress indicator below 40%), the scenario for Ethiopia is expected to shift dramatically in the coming years, considering the increasing trend in historical water stress records. Our water stress model projects that by 2050, Ethiopia’s median water stress will likely range between 300% to 500% under all considered scenarios (refer to Figure 13 in the Appendix C). This implies that Ethiopia would need to secure an additional 3-5 times the current amount of water resources, a situation poised to significantly impact food security risk and, by extension, food production activities. Employing the water stress penalization perspectives from Section 3.3, we gain a more accurate view of Ethiopia’s future food security risk. Notably, even under the NC perspective, a considerable increase in food security risk is anticipated in the forthcoming years across all SSP-RCP scenarios due to the predicted water scarcity hazard. This trend holds true under all perspectives (NC, LC, and VC), indicating a substantial food security risk by 2050, particularly notable in the mid-range SSP2-4.5 scenario.
4.3 Food security risk across SSP-RCPs
While obtaining food security risk estimates within each SSP scenario is insightful, a key limitation of this approach is its lack of robustness to uncertainties regarding which scenario will actually materialize. To address this, and aid policymakers in making informed decisions that consider all possible scenarios, we implement the robust risk assessment approach discussed in Section 3.3. Illustrating this with a toy example, we consider three hypothetical types of policymakers (or experts [^28] For insights into incorporating expert opinions, see Sutherland and Burgman (2015)): (a) one characterized by complete ignorance, hence assigning equal weight to all possible outcomes; (b) an optimist, who assigns higher probabilities to scenarios more favorable for the environment; and (c) a pessimist, leaning towards scenarios less favorable for the environment. These types are detailed in Table 6.
@@ C. Pessimistic 1/100 1/25 1/10 1/5 3/20 1/2
Figure 9 presents the estimates of the food security index across SSP-RCP scenarios, applying the NC, LC, and VC rules for selecting $\gamma$. For Egypt, the estimations under all water stress perspectives align closely, showing a consistent trend across different expert opinions. This trend indicates an increasing concave-type risk in food security up to the year 2040, followed by a slight decline. This pattern is largely attributed to Egypt’s high vulnerability to water scarcity. In the case of Ethiopia, the food security risk estimates also exhibit homogeneity across all perspectives and expert types, revealing an almost exponential increase in food security risk over time. This robust scenario aggregation approach suggests that both countries are facing significant food security challenges, with Ethiopia’s situation predicted to deteriorate more rapidly unless preemptive actions are implemented. Importantly, these estimates across scenarios for both countries demonstrate their robustness, offering consistent predictions that appear relatively unaffected by variations in weights assigned to different SSP-RCP scenarios (reflecting various types of expert opinions).
@@ 5 Conclusions
In this work, we proposed a general methodology for creating probabilistic socioeconomic scenarios aligned with the SSP-RCP framework and for assessing food security risk. This methodology stands out for its emphasis on probabilistic scenarios, which effectively represent the inherent uncertainty and are more suitable for projecting key socioeconomic quantities into the future. As an application, we focused on food security risk, providing a plausible index for assessment and a quantitative evaluation framework per scenario, along with future projections. Acknowledging the uncertainty in which scenario materializes, we also introduced a robust framework for creating projections across scenarios using the concept of convex risk measures in the presence of model uncertainty.
The application of this approach to Egypt and Ethiopia in the upper Nile region reveals diverging food security risks. While Egypt is not expected to face significant food security risk over the next 30 years under any SSP-RCP trajectory, Ethiopia is predicted to experience an exponential increase in water stress, potentially requiring 3-5 times its current water resources by 2050. This alarming projection should urge Ethiopian policymakers to design sustainable water use policies to prevent severe food security issues. Our analysis remains robust across different hypotheses about scenario probabilities, emphasizing the urgent need for Ethiopian policy-makers to explore pathways for food system transformation, as suggested by WHO et al. (2021).
However, the various models employed in this paper are subject to certain assumptions and limitations. The Raftery’s population model and the MaGE, which form the basis of our probabilistic scenario modeling, do not explicitly consider population control policies or climate effects. These are introduced exogenously through RCP scenarios. Future extensions could explore the explicit integration of these effects. Additionally, the data used for calibrating the food capacity model only cover the period up to 2018, excluding recent phenomena like pandemics or political events. Techniques from extreme value theory could be employed to assess the impacts of these shocks.
Finally, while a more detailed food security model incorporating local food production infrastructure, dietary patterns, etc., could be developed, the lack of reliable data for many countries necessitates a more aggregate approach. Thus, our proposed methodology and the resulting food security risk index rely on publicly available, accessible, and reliable data. This approach offers a balanced perspective, providing valuable insights for policymakers while acknowledging the constraints and potential areas for future research.